Problem 13-9 Calculating the WACC Hero Manufacturing has 8.3 million shares of common stock outstanding. The current share price is $79 and the book value per share is $4. The company also has two bond issues outstanding. The first bond issue has a face value of $85 million, a coupon rate of 6.4 percent and sells for 108.1 percent of par. The second issue has a face value of $65.8 million, a coupon rate of 7.6 percent and sells for 111.9 percent of par. The first issue matures in 8 years, the second in 26 years. Suppose the company’s stock has a beta of 1.2. The risk-free rate is 3.6 percent and the market risk premium is 6.7 percent. Assume that the overall cost of debt is the weighted average implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 25 percent.

What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

WAAC ??? %
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=79*8300000
=655700000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*85000*1.081
=91885000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*65800*1.119
=73630200
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=655700000+91885000+73630200
=821215200
Weight of equity = MV of Equity/MV of firm
Weight of equity = 655700000/821215200
W(E)=0.7985
Weight of debt = MV of Bond/MV of firm
Weight of debt = 165515200/821215200
W(D)=0.2015
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 3.6 + 1.2 * (6.7)
Cost of equity% = 11.64
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =8x2
1081 =∑ [(6.4*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^8x2
                   k=1
YTM1 = 5.1517571738
Bond2
                  K = Nx2
Bond Price   =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                   K =26x2
1119 =∑ [(7.6*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^26x2
                    k=1
YTM2 = 6.63
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=5.1517571738*(91885000)/(91885000+73630200)+6.63*(91885000)/(91885000+73630200)
Firm cost of debt=5.81%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.81*(1-0.25)
= 4.3575
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=4.36*0.2015+11.64*0.7985
WACC =10.173%